Suppose you randomly draw a single card from a standard deck of cards.
a. what is the probability you will draw a six of diamonds?
b. what is the probability you will draw a club?
2. Suppose you randomly draw two cards from a standard deck of playing cards. What is the probability of of drawing a four of clubs and then drawing a seven of hearts?

Answer by math_helper(2461)

a. There are 52 cards in the deck, and the 6 of diamonds is just one of them. So Pr{drawing 6 diamonds} = 1/52

b. There are 13 clubs in a deck of cards, so Pr{drawing a club} is 13/52 = 1/4

c. If you randomly chose a card, you have 1/52 chance of that being the exact card you want. If you then draw a 2nd card there is a 1/51 chance of it being the 2nd card you hoped to see. I assume no replacement. Thus, Pr{draw 4 of clubs followed by draw 7 of hearts} = (1/52)(1/51) = 1/2652.